OpenStudy (anonymous):

Is there an inverse for the function f(x)=6x^3+3, or is it not a one-to-one function?

OpenStudy (helder_edwin):

it is a one-to-one function and it has an inverse

OpenStudy (anonymous):

well, I'm taking a quiz, and these are my choices @helder_edwin A) f^-1(x)=cubed root of x+3/6 B) f^-1(x)=cubed root of x-3/6 C) Not a one-to-one function D) cubed root of (x/6)-3

OpenStudy (anonymous):

If it is a one-to-one function, then it must be B. Right?

OpenStudy (helder_edwin):

u have to solve for x the following equation \[\large y=6x^3+3 \]

OpenStudy (helder_edwin):

\[\large y-3=6x^3 \] \[\large \frac{y-3}{6}=x^3 \] \[\large x=\sqrt[3]{\frac{y-3}{6}} \]

OpenStudy (anonymous):

So that's B. Correct?

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

Yay! Thanks so much! I made a hundred on the quiz. God bless you! :D

OpenStudy (helder_edwin):

u r welcome